### v1

# Autonomy Architectures for Safe Planning in Unknown Environments Under Budget Constraints

Abstract
Mission planning can often be formulated as a constrained control problem under multiple path constraints (i.e., safety constraints) and budget constraints (i.e., resource expenditure constraints). In a priori unknown environments, verifying that an offline solution will satisfy the constraints for all time can be difficult, if not impossible. Our contributions are as follows: 1) We propose an online method, building on our previous work gatekeeper, to guarantee safety and satisfy budget constraints of the system trajectory at all times throughout a mission. 2) Next, we prove that our algorithm is recursively feasible and correct. 3) Finally, instead of using a heuristically designed backup controller, we propose a sampling-based method to construct backup trajectories that both minimize resource expenditure and reach budget renewal sets, in which path constraints are satisfied and the constrained resources are renewed. We demonstrate our approach in simulation with a fixed-wing UAV in a GNSS-denied environment with a budget constraint on localization error that can be renewed at visual landmarks. [Code]
I Introduction
Robotic and autonomous systems often operate with limited knowledge of their environment and limited sensing capabilities, which makes designing missions with strict safety requirements challenging. In addition, these systems have a limited budget on resources that are expended throughout a mission. In this paper, we treat safety requirements in the form of a set of states in which the robot must always remain, called the safe set. Specifically, we consider the case where the safe set is not fully known a priori and must be built online with information from the sensor measurements. Online methods require guarantees on recursive feasibility to remain safe if planning methods fail to find a trajectory.
While full trajectory optimization methods exist to handle constrained problems [1], they are often non convex and computationally intractable to solve online. Path planning techniques [2] can generate trajectories quickly based on approximated dynamics, but the solutions may not be dynamically realizable and safety guarantees cannot be guaranteed.
Constraints on finite resources have been studied under various names, including integral constraints [3, 4], budget constraints [5, 6], and cost constraints [7, 8]. Examples include constraints on battery charge [9], localization error [10], or time spent visible to an enemy observer [11]. We consider a generalized case of the budget-constrained planning problem where resources can be renewed at specific regions in the state space.
Various methods to enforce budget constraints have been designed in previous works, ranging from solutions for specific systems and types of missions to general solutions solved offline. An example of a specific application of localization error constrained planning for UAVs is in [8], in which a greedy multi-objective motion planning strategy is used to navigate with little knowledge of the environment, making use of cellular signals to localize the robot.
In [12], a control barrier function (CBF) is used to enforce battery charge constraints, but it assumes a simple, differentiable battery model and is limited to the class of systems with single-integrator dynamics. The methods in [5, 3] solve full optimal control problems with budget constraints and resets in discrete and continuous state spaces, but these approaches rely on large-scale graph searches or solving viscosity equations derived from dynamic programming, which are computationally expensive and better suited for offline planning in known environments.
We note that none of these solutions explicitly handle the combination of general nonlinear dynamics, limited environmental knowledge, and non-trivial budget expenditure dynamics, all while guaranteeing safety. In addition, some of these methods involve mission- and constraint-specific design of the planning modules, which reduces modularity and adaptability of the system to additional types of constraints and missions. To that end, we aim to build a layered control architecture with modular components [13].
Our proposed solution to the problem of ensuring safety and budget constraint guarantees for nonlinear systems in unknown environments builds on our previous work (gatekeeper [14]) to incorporate budget constraints and provides a method for constructing backup trajectories. gatekeeper is a modular framework that ensures safety by recursively constructing a candidate trajectory consisting of a nominal trajectory, followed up to some switching time, and a backup trajectory that guarantees constraint satisfaction thereafter. Recently, gatekeeper has been applied in the eware filter which is designed specifically for battery life constraints, where energy is renewed at a charging station [9, 15]. Our solution further generalizes gatekeeper to enforce a budget constraint on a resource which is expended during the mission and can be renewed in designated regions.
In summary, this paper has three main contributions:
  • We propose an architecture to guarantee safety and satisfy budget constraints while following a nominal trajectory in unknown environments, based on gatekeeper.
  • We prove recursive feasibility and correctness, ensuring that system trajectories satisfy all constraints at all times.
  • We propose a sampling-based method called ReRoot to construct a backup policy that minimizes resource expenditure while ensuring the system reaches budget renewal sets, sets where safety constraints hold and resources are replenished.
The paper is organized as follows. In Section II, we introduce the problem formulation. Section III describes the method overview of gatekeeper and ReRoot. Section IV discusses a simulation case study of a UAV in a GNSS-denied environment navigating to a goal with safety and budget constraints related to visual odometry.
II Problem Formulation
II-A System Modeling
We represent the robot motion by the nonlinear dynamics:
$\dot{x}=f(x,u)$ (1)
where $x\in\mathcal{X}\subset\mathbb{R}^{n}$ is the state and $u\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control input, and $f:\mathcal{X}\times\mathcal{U}\to\mathbb{R}^{n}$ a locally Lipschitz function. Under a (locally Lipschitz) feedback controller $\pi:\mathcal{X}\to\mathcal{U}$ and an initial condition $x(t_{0})=x_{0}\in\mathcal{X}$, the closed-loop system dynamics read:
$\dot{x}=f(x,\pi(x)),~{}x(t_{0})=x_{0}.$ (2)
We consider a resource that is subject to a budget constraint, and which can be renewed in a subset $\mathcal{R}$ of the state space $\mathcal{X}$. The budget state $b$ is governed by the following hybrid dynamical system:
$\begin{cases}\dot{b}=L(x,u),\quad&x\notin\mathcal{R}\\ b^{+}=b_{\textup{reset}},\quad&x\in\mathcal{R}\end{cases}$ (3)
with initial condition
$b(t_{0})=\begin{cases}b_{0},\quad&x\notin\mathcal{R}\\ b_{\textup{reset}},\quad&x\in\mathcal{R},\end{cases}$ (4)
where $L:\mathcal{X}\times\mathcal{U}\to\mathbb{R}_{\geq 0}$ is piecewise continuous and $b^{+}=b_{\textup{reset}}\geq 0$ is the reset value of the budget state after the jump. We assume that $L$ is known, but $\mathcal{R}$ is not assumed to be fully known.
II-B Set Representation of the Environment
We aim to satisfy the following constraints:
$x(t)\in\mathcal{S},$ $\quad\forall t\geq t_{0}$ (5)
$b(t)\leq B,$ $\quad\forall t\geq t_{0}$ (6)
where $\mathcal{S}\subset\mathcal{X}$ is the set of states that satisfy mission safety requirements and $B$ is the budget constraint that must not be exceeded. We assume that $\mathcal{S}$ has a nonempty interior and that $B>b_{\textup{reset}}$.
We assume $\mathcal{S}$ is not fully known, but can be sensed online from sensor measurements. We introduce the known subset of the safe set at time $t_{k}$, $\mathcal{F}_{k}\subset\mathcal{S}$, where $k\in\mathbb{N}$. Similarly, we denote the known renewal set as $\mathcal{R}_{k}\subset\mathcal{R}\subset\mathcal{S}$, where we have assumed the renewal set is safe.
At time $t_{k+1}$, new information is acquired by the robot to update $\mathcal{F}_{k}$ and $\mathcal{R}_{k}$ such that the following relations hold:
$\mathcal{F}_{k}\subset\mathcal{F}_{k+1}\quad\forall k\in\mathbb{N}$ (7)
$\mathcal{R}_{k}\subset\mathcal{R}_{k+1},~{}\forall k\in\mathbb{N}.$ (8)
Moreover, $\mathcal{R}_{k}$ can be composed of $N_{C}\geq 1$ disjoint, compact subsets. At time $t_{k}$, we denote the $i$-th subset as $\mathcal{R}^{i}_{k}$ for $i\in\{1,\dotsc,N_{C}\}$ such that $\mathcal{R}_{k}=\cup_{i=1}^{N_{C}}\mathcal{R}^{i}_{k}$.
Our method relies on concepts in set invariance to guarantee constraint satisfaction [16].
Definition 1 (Controlled Invariant Set).
A set $\mathcal{C}\subset\mathcal{X}$ is controlled invariant for the system (1), if there exists a controller $\pi:\mathcal{X}\to\mathcal{U}$ that assures the existence of a unique solution to the closed-loop system (2) such that $\mathcal{C}$ is positively invariant for the closed-loop system (2), i.e., $\forall x(t_{0})\in\mathcal{C},x(t)\in\mathcal{C},\forall t\geq t_{0}.$
Definition 2 (Backup Set).
A set $\mathcal{B}\subset\mathcal{X}$ is a backup set if there exists a controller $\pi:\mathcal{X}\to\mathcal{U}$ such that for the closed-loop system (2), $\forall x(t_{0})$ in a neighborhood $\mathcal{N}\subset\mathcal{X}$, there exists a finite time $0\leq T_{B}(x(t_{0}))<\infty$ such that $\mathcal{B}$ is reachable at $t_{0}+T_{B}$ and controlled invariant thereafter, i.e., $x(t)\in\mathcal{B},~{}\forall t\geq t_{0}+T_{B}$.
Assumption 1.
We assume that the budget renewal set $\mathcal{R}_{k}$ is a backup set for the closed loop system (2) for all $k\in\mathbb{N}$. Then, from (3), we obtain:
$x(t)\in\mathcal{R}_{k}\implies b(t)=b_{\textup{reset}}\leq B,\quad\forall t\geq t_{k}.$ (9)
II-C Nominal Trajectory
Definition 3 (Trajectory).
A trajectory with horizon $T_{H}$ is a piecewise continuous function $p:\mathcal{T}\to\mathcal{X}$ defined on $\mathcal{T}=[t_{0},t_{0}+T_{H}]\subset\mathbb{R}_{\geq 0}$. A trajectory $p$ is dynamically feasible w.r.t. (1) if there exists a piecewise continuous control $u:\mathcal{T}\to\mathcal{U}$ s.t.
[…]
Assumption 2.
We assume there exists a feedback controller $\pi^{T}:\mathcal{X}\times\mathcal{X}\to\mathcal{U}$, where $u(t)=\pi^{T}(x(t),p(t))$, which tracks a given trajectory $p$ by minimizing the error $\left\|x(t)-p(t)\right\|$ for $t\geq t_{0}$. We refer to this policy as the tracking controller.
We assume there exists a nominal planning module that computes a nominal trajectory that achieves the mission objectives. We denote the nominal trajectory from time $t_{k}$ as $p_{k}^{\textup{nom}}:[t_{k},t_{k}+T_{H}]\to\mathcal{X}$. The nominal trajectory is not required to be dynamically feasible w.r.t. (1). Formally, we state the following assumption:
Assumption 3.
For each $k\in\mathbb{N}$, a nominal trajectory $p_{k}^{\textup{nom}}:[t_{k},t_{k}+T_{H}]\to\mathcal{X}$ is known.
Figure 1: Block diagram of our proposed layered autonomy architecture, with our module gatekeeper + ReRoot highlighted in the dashed line.
Figure 2: Snapshots of gatekeeper with ReRoot trees rooted at budget renewal sets. The trees are grown in the free set $\mathcal{F}_{k}$, with the path from any node to the root being optimal w.r.t. to the budget cost. The robot discovers new budget renewal sets as it uncovers more of the unknown space.
II-D Problem Statement
With the above preliminaries, we state the main problem:
Problem 1.
Given the known safe set $\mathcal{F}_{k}$ and budget renewal set $\mathcal{R}_{k}$ updated at each $k\in\mathbb{N}$, design an architecture such that the closed-loop system (2) tracks a nominal trajectory $p^{\textup{nom}}_{k}$ while guaranteeing all safety and budget constraints are satisfied, i.e., $x(t)\in\mathcal{S}$ and $b(t)\leq B$ for all $t\geq t_{0}$.
III Method Overview
To address Problem 1, we propose an architecture built around gatekeeper, a safety filter, and ReRoot, a backup policy generation algorithm. In particular, we develop a solution that offers two contributions compared to the earlier version of the gatekeeper framework: 1) it accounts for budget constraints, and 2) we provide a constructive method to generate a backup policy that aims to maximize the time spent tracking the desired trajectories. See Figure 2 for a block diagram showing where gatekeeper fits into the general autonomy stack and Table I for a summary of the nomenclature in this paper.
TABLE I: Nomenclature.
Symbol Definition
$t_{k}$ Start time of gatekeeper iteration $k$
$t_{kS}$ Trajectory switching time $t_{kS}=t_{k}+T_{S}$
$t_{kB}$ Backup trajectory time $t_{kB}=t_{kS}+T_{B}$
$\mathcal{X}\subset\mathbb{R}^{n}$ State space
$\mathcal{U}\subset\mathbb{R}^{m}$ Control input space
$\mathcal{S}\subset\mathcal{X}$ Safe set at time
$\mathcal{F}_{k}\subset\mathcal{S}$ Known safe set at time $t_{k}$
$\mathcal{R}_{k}\subset\mathcal{S}$ Budget renewal set at time $t_{k}$
$\mathcal{G}=(\mathcal{V},\mathcal{E})$ ReRoot graph $\mathcal{G}$ with nodes $\mathcal{V}\subset\mathcal{X}$ and edges $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$.
$p_{k}^{\textup{nom}}$ $k$-th nominal trajectory on $[t_{k},t_{k}+T_{H}]$
$p_{k}^{\textup{bck},T_{S}}$ backup trajectory starting at time $t_{k}+T_{S}$
$p_{k}^{\textup{can},T_{S}}$ candidate trajectory at time $t_{k}$ with switching time $T_{S}$
$p_{k}^{\textup{com}}$ $k$-th committed trajectory on $[t_{k},t_{kB}]$
First, we summarize the key steps of our method. In the following subsections, we provide mathematical definitions of each type of trajectory along with a proof of correctness. gatekeeper is an online safety verification module that recursively filters the nominal planner output by constructing a trajectory that is guaranteed to satisfy all mission constraints [14]. At the $k$-th iteration of our method, where $k\in\mathbb{N}\setminus 0$, the following steps occur:
  • For all $T_{S}\in[0,T_{H}]$, propagate the nominal trajectory $p_{k}^{\textup{nom}}$ (black line in Figure 2) on the interval $[t_{k},t_{k}+T_{S}]$, then generate a backup trajectory $p^{\textup{bck},T_{S}}_{k}$ that reaches $\mathcal{R}_{k}$ with ReRoot, which forms a set of candidate trajectories.
  • For each candidate trajectory, compute the budget state time history $b(t)$ with (3) and verify that constraints (5) and (6) are satisfied, forming a set of valid trajectories. In Figure 2 when $k=2$, candidate trajectories that enter the red unsafe area are deemed invalid.
  • Select the valid trajectory that maximizes $T_{S}$ as the committed trajectory $p_{k}^{\textup{com}}$. In Figure 2, the committed trajectory is composed of a red nominal portion and a green backup portion, which is generated from ReRoot and will be detailed in the following subsections. If there are no valid trajectories, select the $(k-1)$-th committed trajectory $p^{\textup{com}}_{k-1}$. This step guarantees recursive feasibility.
Overall, Figure 2 depicts the gatekeeper trajectories and ReRoot growth as the robot expands the known budget renewal subsets $\mathcal{R}_{k}^{i}$ during the mission.
Our goal is to formally prove that all committed trajectories are guaranteed to satisfy all constraints for all time. To do so, we need to first introduce the definitions below.
III-A Guaranteed Constraint Satisfaction
We formally introduce the notion of a backup trajectory.
Definition 4 (Backup Trajectory).
A trajectory $p:\mathcal{T}\to\mathcal{X}$ over the interval $\mathcal{T}=[t_{k}+T_{S},t_{k}+T_{S}+T_{B}]$ is a backup trajectory $p^{\textup{bck},T_{S}}_{k}:\mathcal{T}\to\mathcal{X}$ to a set $\mathcal{B}\subset\mathcal{X}$ over the interval if the resulting control policy $u=\pi^{T}(x,p^{\textup{bck},T_{S}}_{k})$ renders $\mathcal{B}$ a backup set by Definition 2 at time $t_{kB}=t_{k}+T_{S}+T_{B}$.
Note that the backup trajectory is not required to satisfy constraints (5) and (6) – the validation step of the candidate trajectories will filter out any unsafe backup trajectories, which will be detailed below. As gatekeeper maximizes the time spent tracking the nominal trajectory, the construction of the backup trajectory is key to avoid early switching. In the next subsection, we detail our method for constructing backup trajectories with ReRoot.
The candidate trajectories are constructed from the nominal trajectory and some backup trajectories.
Definition 5 (Candidate Trajectory).
Given a switching time $T_{S}\in[0,T_{H}]$, the solution to the initial value problem
$\dot{p}$ $=f(p,u(t)),~{}p(t_{k})=x_{k},$ (11)
$u(t)$ =$=\begin{cases}\pi^{T}(p(t),p_{k}^{\textup{nom}}(t)),&\quad t\in[t_{k},t_{k}+T_{S})\\ \pi^{T}(p(t),p_{k}^{\textup{bck},T_{S}}(t)),&\quad t\geq t_{k}+T_{S}\end{cases}$ (12)
is the candidate trajectory $p_{k}^{\textup{can},T_{S}}$ defined on $[t_{k},\infty)$.
A candidate trajectory is valid if it satisfies all constraints at every point along the trajectory and reaches a budget renewal set in finite time.
Definition 6 (Valid Trajectory).
A candidate trajectory is valid if it remains in the known safe set:
$p_{k}^{\textup{can},T_{S}}(t)\in\mathcal{F}_{k}\quad\forall t\in[t_{k},t_{kB}],$ (13)
the solution to (3) remains below the budget $B$:
$b(t)\leq B\quad\forall t\in[t_{k},t_{kB}],$ (14)
and the trajectory reaches $\mathcal{R}_{k}$:
$p_{k}^{\textup{can},T_{S}}(t_{kB})\in\mathcal{R}_{k}.$ (15)
Finally, the committed trajectory is chosen to be the valid trajectory that maximizes $T_{S}$. If no committed trajectory can be found, the previously committed trajectory is followed.
Definition 7 (Committed Trajectory).
At the $k$-th iteration, define the set of valid candidate trajectories parameterized by $T_{S}$ as
[…]
where $p_{k}^{\textup{can},T_{S}}$ is as defined in (11).
If $\mathcal{I}_{k}\neq\emptyset$, let $T_{S}^{*}=\max\mathcal{I}_{k}$. The committed trajectory is
$p_{k}^{\textup{com}}(t)=p_{k}^{\textup{can},T_{S}^{*}}(t),\quad t\in[t_{k},\infty).$ (17)
If $\mathcal{I}_{k}=\emptyset$, the committed trajectory is
$p_{k}^{\textup{com}}(t)=p_{k-1}^{\textup{com}}(t),\quad t\in[t_{k},\infty).$ (18)
Finally, we can state the following theorem that gatekeeper guarantees constraint satisfaction for all time.
Theorem 1.
Suppose $p_{0}^{\textup{can},T_{S}}$ is a dynamically feasible candidate trajectory defined on $[t_{0},\infty)$ that is valid by Definition 6 for some $T_{S}\geq 0$. If $p_{k}^{\textup{com}}$ is determined via Definition 7, then
$p_{k}^{\textup{com}}(t)\in\mathcal{S},\quad\forall t\in[t_{k},\infty),$ (19)
$b(t)\leq B,\quad\forall t\in[t_{k},\infty).$ (20)
Furthermore, if $x(t_{0})=p_{0}^{\textup{com}}(t_{0})$, and the system (1) is driven by $u(t)=\pi^{T}_{k}(x(t),p_{k}^{\textup{com}}(t))$, for $t\in[t_{k},t_{k+1})$, then the closed-loop dynamics (2) satisfy $x(t)\in\mathcal{S}$ and $b(t)\leq B$ for all $t\geq t_{0}$.
Proof.
The proof is by induction and is based on Theorem 1 in [14].
Base Case: Let $k$ = 0. Since $p^{\textup{can},T_{S}}_{0}$ is valid, it is committed as $p_{0}^{\textup{com}}=p^{\textup{can},T_{S}}_{0}$. Then for the committed trajectory,
$p_{0}^{\textup{com}}(t)\in\begin{cases}\mathcal{F}_{0},\quad t\in[t_{0},t_{0,B})\\ \mathcal{R}_{0},\quad t\geq t_{0,B}\end{cases}$
$\implies$ $p_{0}^{\textup{com}}(t)\in\begin{cases}\mathcal{S},\quad t\in[t_{0},t_{0,B})\\ \mathcal{S},\quad t\geq t_{0,B}\end{cases}$
$\iff$ $p_{0}^{\textup{com}}(t)\in\mathcal{S},\quad\forall t\geq t_{0}.$
where $t_{0,B}=t_{0}+T_{S}+T_{B}$. We used the fact that $\mathcal{R}_{0}$ is a backup set by Definition 2 and is thus controlled invariant. Hence, $p_{0}^{\textup{com}}$ satisfies all constraints for all time. Moreover, since $p_{0}^{\textup{com}}$ is dynamically feasible, $x(t)=p_{0}^{\textup{com}}(t)\quad\forall t\in[t_{0},t_{1})$. Then for the budget state solution to (3) with $x(t)$ and the corresponding $u(t)$,
$b(t)\begin{cases}\leq B,\quad t\in[t_{0},t_{0B})\\ =b_{\textup{reset}}\leq B,\quad t\geq t_{0B}\end{cases}$
$\iff b(t)\leq B,\quad\forall t\geq t_{0}.$
Induction Step: Now suppose the claim in Theorem 1 is true for some $k\in\mathbb{N}$. To show the claim holds for $k+1$, consider the two possible definitions for $p_{k+1}^{\textup{com}}$ from Definition 7:
Case 1: If $\mathcal{I}_{k+1}\neq\emptyset$, then
$p_{k+1}^{\textup{com}}(t)$ $=p_{k+1}^{\textup{can},T_{S}^{*}}(t)~{}\forall t\geq t_{k+1}$
$\in\begin{cases}\mathcal{F}_{k+1},\quad t\in[t_{k+1},t_{k+1,SB})\\ \mathcal{R}_{k+1},\quad t\geq t_{k+1,SB}\end{cases}$
$\in\mathcal{S},\quad\forall t\geq t_{k+1}.$
Case 2: If $\mathcal{I}_{k+1}=\emptyset$, the committed trajectory is unchanged
$p_{k+1}^{\textup{com}}(t)=p_{k}^{\textup{com}}(t)\in\mathcal{S},\quad\forall t\geq t_{k+1}.$
Then since $p_{k+1}^{\textup{com}}$ is dynamically feasible by construction, $x(t)=p_{k+1}^{\textup{com}}(t)\quad\forall t\in[t_{k+1},t_{k+2})$. Therefore, $x(t)=p_{k}^{\textup{com}}(t)\in\mathcal{S}~{}\forall t\in[t_{k},t_{k+1}),~{}\forall k\in\mathbb{N}\implies x(t)\in\mathcal{S}~{}\forall t\geq t_{0}$. Moreover, the solution to (3) with the corresponding $x(t)$ and $u(t)$ satisfies $b(t)\leq B~{}\forall t\geq t_{0}$. ∎
We have shown that using the gatekeeper architecture with a backup trajectory that reaches $\mathcal{R}_{k}$, the mission requirements are satisfied for all time. Next, we introduce a method for constructing backup trajectories and selecting one that minimizes the budget state $b$.
III-B Backup Trajectory Construction
Our method relies on posing a trajectory optimization problem and solving it with a sampling-based algorithm called ReRoot.
III-B1 Trajectory Optimization Problem
Consider a valid candidate trajectory $p(t)$ and the corresponding budget state prediction $b(t)$. Let the pre-jump budget state be the one-sided limit as $t$ approaches $t_{kB}$ from below:
$b^{-}=\lim_{t\nearrow t_{kB}}b(t),$ (21)
where $t_{kB}=t_{k}+T_{S}+T_{B}$ is the finite time when $p$ reaches $\mathcal{R}_{k}$.
For any switch time $T_{S}$, the ideal backup trajectory is the one that minimizes the budget expenditure. Given a known budget renewal set $\mathcal{R}_{k}$ at time $t_{k}$, design a backup trajectory planner that computes a dynamically feasible trajectory $p^{\textup{bck},T_{S}}_{k}$ starting at $p_{k}^{\textup{can},T_{S}}(t_{kS})$ and ending in $\mathcal{R}_{k}$ while remaining in the known safe set $\mathcal{F}_{k}$ with minimal budget expenditure along the backup trajectory:
$\min_{\begin{subarray}{c}p(\cdot),b(\cdot),u(\cdot)\end{subarray}}\quad$ $b^{-}$ (22a)
s.t. $\dot{p}(t)=f(x(t),u(t))$ (22b)
[…]
$p(t_{kS})=p_{k}^{\textup{can},T_{S}}(t_{kS})$ (22d)
$b(t_{kS})=b_{k}^{\textup{can},T_{S}}(t_{kS})$ (22e)
$x(t)\in\mathcal{F}_{k}\quad\forall t\in[t_{kS},t_{kB}]$ (22f)
$u(t)\in\mathcal{U}\quad\forall t\in[t_{kS},t_{kB}]$ (22g)
$p(t_{kB})\in\mathcal{R}_{k}$ (22h)
Here, (22b) and (22c) are the robot and budget dynamics, (22d) and (22e) are the trajectory and budget initial conditions based on the candidate trajectory and corresponding propagated budget state $b_{k}^{\textup{can},T_{S}}$, (22f) is the safety constraint, (22g) is the input constraint, and (22h) is the terminal constraint on the trajectory $p$.
We will solve (22) for various initial conditions with a sampling-based method detailed below.
III-B2 ReRoot (Reverse Rooted Forest)
We build a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with a set of nodes $\mathcal{V}\subset\mathcal{X}$ and a set of edges $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$, in a similar manner to the standard RRT* algorithm [17]. The key differences between ReRoot and RRT* are as follows:
  • The graph $\mathcal{G}$ is a rooted forest, i.e., a union of disjoint rooted trees, where the roots are in the renewal set $\mathcal{R}_{k}$. Furthermore, in the rewiring step, it is possible for a node (and its children) to switch to a different tree with a different root.
  • We grow the forest in reverse such that the roots are backup “goal” states in $\mathcal{R}_{k}$ and the tree extends to cover $\mathcal{F}_{k}$, which results in optimal paths from all states in $\mathcal{F}_{k}$ that end in $\mathcal{R}_{k}$. In RRT* the initial state is the root and the tree is grown until it reaches the goal state.
  • Instead of minimizing path length, the paths minimize the budget state at the end of the backup trajectory $b^{\textup{bck}}$. Since the budget dynamics only depend on the state $x$ and control $u$, the cost of the path from a node to the root is the sum of the incremental costs along the path.
Despite these differences in the interpretation of the nodes and paths, the core RRT* algorithm is unchanged and provides asymptotic optimality of the cost of all paths as $|\mathcal{V}|\to\infty$.
Next, we describe the process of extracting a backup trajectory from the ReRoot graph.
III-B3 Backup Trajectory Generation
Denote $d:\mathcal{V}\to\mathbb{R}_{\geq 0}$ as the depth of a node, which is the number of edges between $v$ and the root of its tree. A root node has depth $d=0$. For $v\in\mathcal{V}$, define
$\mathcal{W}(v)=\{v,\mathcal{P}(v),\mathcal{P}^{2}(v),\dotsc,\mathcal{P}^{d}(v)\}$ (23)
as the set of waypoints from $v$ to the root of its tree, where $\mathcal{P}:\mathcal{V}\to\mathcal{V}$ is the parent node of $v$.
Then, for each $v\in\mathcal{V}$, let $p_{v}$ be the closed loop trajectory from tracking the waypoints $\mathcal{W}(v)$:
$\dot{p}_{v}=f(p_{v},\pi^{T}(\mathcal{W}(v)),~{}p_{v}(t_{kS})=p^{\textup{can},T_{S}}_{k}(t_{kS}).$ (24)
Additionally, let $b_{v}$ be the budget state expenditure over the trajectory $p_{v}$ by solving (3).
Finally, the trajectory with minimal $b_{v}$ that is safe for all time is chosen as the backup trajectory:
$p^{\textup{bck},T_{S}}_{k}=\operatorname*{\arg\!\min}_{p_{v}}\{b_{v}~{}|~{}p_{v}(t)\in\mathcal{F}_{k}~{}\forall t\geq t_{kS}\}.$ (25)
Note that by construction, $p^{\textup{bck},T_{S}}_{k}$ is dynamically feasible and guaranteed to remain in the known safe set $\mathcal{F}_{k}$. The resulting candidate trajectory $p^{\textup{can},T_{S}}_{k}$ is the solution to (11). The validity of the candidate trajectory is given from Definition 6, i.e., checking that $b(t)\leq B$ by forward propagating (3).
III-C Summary
In summary, we have provided a solution to 1 with a layered architecture built on the gatekeeper trajectory filter and the ReRoot backup policy. To guarantee that safety and budget constraints are satisfied for all time as the nominal trajectory is tracked, a backup trajectory is planned and stitched with the largest portion of the nominal trajectory such that the resulting committed trajectory satisfies all constraints and reaches a budget renewal set. We then provided a method for constructing backup trajectories called ReRoot that minimizes the expenditure of the budget state to allow the nominal trajectory to “consume” more of the budget. ReRoot is a reverse rooted forest in which branches represent paths of minimal budget cost that start from nodes in the known safe space and reach a budget renewal set.
IV Simulation Case Study
IV-A Simulation Setup
We evaluate our method in a simulation case study of a fixed-wing UAV flying in a GNSS-denied environment, using visual-inertial odometry to estimate its position. The horizontal planar dynamics are modeled as a Dubins vehicle:
$\dot{r}_{N}$ $=V\cos\psi$ (26a)
$\dot{r}_{E}$ $=V\sin\psi$ (26b)
$\dot{\psi}$ $=u$ (26c)
where $[r_{N},r_{E}]^{\intercal}\in\mathbb{R}^{2}$ is the horizontal position in a North-East-Down (NED) coordinate frame, $\psi\in[0,2\pi)$ is the heading, $V>0$ is a fixed velocity and $u$ is the input turn rate. $u$ is constrained by the minimum turn radius of the UAV, $\rho>0$, such that $u\in\mathcal{U}=[-\frac{V}{\rho},\frac{V}{\rho}]$. For simplicity and visualization purposes, we constrain the UAV to fly at a fixed altitude, but our method extends trivially to 3D trajectories, or for more complicated dynamics.
The UAV’s velocity is $V=10$ m/s and minimum turn radius is $\rho=10$ m, which corresponds to a $\sim$45° bank angle. The sensor field of view (FoV) is defined by a radius of 60 m and an angle of 90°. The perception and mapping module runs at 5 Hz.
The objective to reach the randomly chosen goal location within the 250 m $\times$ 200 m mission domain. The mission requirements are as follows:
  • (Safety constraint) Maintain at least $N_{f}=8$ visual features in the FoV between successive camera frames for robust feature tracking.
  • (Budget constraint) Maintain an absolute position error (error bound on position) below $B=9$ m.
Note that the safety constraint is difficult to write in equation form, e.g., $h(x)\leq 0$, and is not differentiable. It is however easy to verify algorithmically. We assume there exist some visual landmarks that reset position error to zero when the UAV can see them. We assume that the position error is proportional to distance traveled, with an error rate of 3%, i.e, 0.03 m of error per 1 m traveled, based on [18].
Translating these requirements b into the format needed for our method, we first write the safe set as
$\mathcal{S}=\{x\in\mathcal{X}~{}|~{}\textup{FeaturesInFoV}(x)\geq N_{f}\textup{ or }x\in\mathcal{R}\},$ (27)
where $\textup{FeaturesInFoV}:\mathcal{X}\to\mathbb{N}$ is the number of visible features at state $x$. We assume no prior knowledge of the feature locations. The UAV maps the features as it flies through the environment and builds $\mathcal{F}_{k}$. The budget state models the growth of localization error as
$\begin{cases}\dot{b}=0.03V,\quad&x\notin\mathcal{R}\\ b^{+}=b_{\textup{reset}}=0,\quad&x\in\mathcal{R}\end{cases}$ (28)
with initial condition
$b(t_{0})=\begin{cases}b_{0},\quad&x\notin\mathcal{R}\\ 0,\quad&x\in\mathcal{R}.\end{cases}$ (29)
We define the budget renewal set to have subsets $\mathcal{R}^{i}$ at the starting location, a mid-point landmark (unknown to the UAV at the mission start time), and around the known goal location. Since the UAV cannot hover at a single location under the dynamics (26), we define $\mathcal{R}$ as a circular counter-clockwise orbit of radius $\rho$ around the landmarks. $\mathcal{R}$ is controlled invariant under the constant control input $u=-\frac{V}{\rho}$ and reachable from nearby states, i.e., a backup set. The definition for $\mathcal{R}^{i}$ is as follows:
$\mathcal{R}^{i}=\Big{\{}x\in\mathcal{X}~{}|~{}(r_{N}-l_{N}^{i})^{2}+(r_{E}-l_{E}^{i})^{2}=\rho^{2},$ (30)
$\psi=\arctan\left(\frac{r_{E}-l_{E}^{i}}{r_{N}-l_{N}^{i}}\right)-\frac{\pi}{2}\Big{\}},$ (31)
where $l_{N}^{i}$ and $l_{E}^{i}$ are the north and east positions of the $i$-th landmark. All backup trajectories must end in a subset $\mathcal{R}^{i}$, which means the UAV can orbit indefinitely while not accumulating localization error.
Figure 3: Top-down view of field environment from the VPAIR database [19]. The white dots are visual odometry features. Note the lack of features in certain areas. The orange circle is the starting location. The yellow circle is the goal location. The white circle is a landmark. The nominal trajectory in black is the path of minimum distance, which becomes unsafe at the red “!”. The magenta line is the omniscient trajectory, i.e., has knowledge of all feature locations, that minimizes distance while satisfying the safety and budget constraints.
Figure 4: (a)-(d) Visualization of the simulated experiment of gatekeeper with ReRoot at various times in the mission. The features seen by the UAV are the white dots. The orange dot is the starting location, the gray dot is a landmark, and the yellow dot is the goal location. The colored thin lines are the branches of the ReRoot forest. The blue line is the UAV’s path up to time $t$. The black line is the unsafe part of the nominal trajectory, and the red line is the nominal component of the committed trajectory, which is replanned from the UAV position when the committed trajectory deviates from the nominal. The green line is the backup component of the committed trajectory and reaches a budget renewal set. (e) Value of the budget state $b$ (absolute position error) over time, which resets whenever the UAV reaches a budget renewal set. The budget never exceeds the maximum allowed value of 9 m. (f) Number of features in the FoV over time, which is never below the minimum $N_{f}=8$. The gray refers to when the UAV is at a landmark.
Figure 5: (a) Simulation result of the same setup as in Figure 4 but with a budget constraint of 5 m of localization error. (b) Despite never reaching the goal, the budget constraint is always satisfied and (c) the safety constraint is also always satisfied. A limit cycle-like behavior is observed, illustrating that gatekeeper can successfully prevent the robot from leaving the safe set when the mission cannot be executed safely.
The simulated environment is a top-down image of a field taken from an airplane as part of the VPAIR database [19] and we create the set of features using Good Features to Track [20], as shown in Figure 3. The mission domain is a 265 m $\times$ 200 m region. This environment presents a challenging setting for visual odometry, as certain regions of the field contain sparse or no discernible features. gatekeeper with ReRoot ensures that the UAV does not enter these unsafe regions. In the simulation with gatekeeper and ReRoot, the UAV has no knowledge of feature locations a priori and maps them online when they enter the FoV. The mid-field landmark is also unknown to the UAV and is discovered during the mission. We further discuss the trajectories in Figure 3 in the results subsection below.
For the nominal planner, we use RRT* in a similar manner to the method in [21], which facilitates replanning from any location in the map and returns a trajectory that is dynamically feasible for the Dubins dynamics (26).
We initialize ReRoot around the starting location to form a tree of safe backups. Then, at 1 Hz, nodes are added based on the newly mapped features to extend the tree into the known safe set. As landmarks are discovered, a root node is placed and nearby nodes automatically connect to the new tree.
IV-B Results and Discussion
The Dubins dynamics in (26) allows fast backup trajectory generation because the edges of ReRoot represent dynamically feasible subpaths. Given two nodes $v_{1},v_{2}\in\mathcal{V}$, the Dubins path of minimum distance is unique and is found algebraically [22]. Hence any list of waypoints $\mathcal{W}(v)$ can be tracked perfectly and the budget state $b(t)$ along the backup trajectory is exactly the node cost of $v$, which is used when building ReRoot to determine the optimal parent of a new node. In the more general case, the trajectories can be computed numerically by solving the differential equations, as in [14].
Figure 3 depicts the nominal trajectory from the start to the goal, which becomes unsafe as there are fewer than $N_{f}=8$ features in the FoV. We include for comparison the omniscient trajectory that satisfies the safety and budget constraints in which the feature locations are known a priori.
The code and animations of the simulation are available here [^1] https://github.com/dcherenson/gatekeeper_budget_constraints. In the simulation, we observe that the UAV reaches the goal while satisfying all constraints in Figure 4. Figure 4a shows that the nominal trajectory across the sparse field is found to be unsafe, so a short backup trajectory is planned to return to its orbit. In Figure 4b, the UAV has not mapped enough of the area around the mid-point landmark to reach it safely and must return to the starting location to reset the localization error. Then, in Figure 4c, the UAV reaches the landmark using the ReRoot tree rooted at the landmark. Finally, in Figure 4d, the UAV is able to reach the goal location with a committed trajectory back to the landmark. Figure 4e shows the position error over time which never exceeds the budget constraint. Figure 4f shows the number of visible features over time, showing that the safety constraint is always satisfied.
The behavior of the UAV trajectory in Figure 4 shows the exploratory effects of gatekeeper. Since following the nominal trajectory to the goal is found to be unsafe, the UAV turns back towards the start. During the turn, it maps more features and is able to extend further into the newly mapped known safe set. Eventually, the budget constraint forces the return to the start to renew the accumulated error. Then, the nominal trajectory can be followed for a longer time as the known safe set has been mapped. This process continues until the UAV reaches the goal while mapping its environment. The shape of the resulting trajectory resembles the omniscient path in Figure 3 with the addition of multiple return trips to landmarks.
We also simulate the mission with a lower, more challenging budget of $B=5$ m of accumulated localization error, which makes the task of reaching the goal while satisfying all constraints infeasible. In Figure 5a, the UAV is unable to reach the landmark or the goal without exceeding the limit. The resulting pattern is caused by the UAV following the nominal trajectory until gatekeeper prevents the UAV from continuing to avoid exceeding the budget, then returning to the home landmark, and the cycle continues. Figure 5b and Figure 5c show that the safety and budget constraints are satisfied, despite this infinite looping behavior. This demonstrates gatekeeper’s ability to ensure constraint satisfaction even when the mission objectives cannot be satisfied safely.
V Conclusion
In this paper, we proposed an architecture centered on gatekeeper, an online safety verification algorithm, to guarantee that safety and budget constraints are enforced throughout a mission. We proved that the method is recursively feasible and guarantees safety for all time. We then introduced ReRoot, a method to construct a backup trajectory that allows for efficient online trajectory generation to minimize the expenditure of the budget state. The algorithm grows multiple rooted trees in reverse starting from budget renewal sets and extending into the known safe set.
We demonstrated our algorithm in simulation with a case study of a fixed-wing UAV in a GNSS-denied environment navigating to a goal location. The safety constraint was to maintain a minimum of 8 visual features in the FoV and the budget constraint was to limit the localization error to below 9 m. The combination of gatekeeper with ReRoot showed that multiple backup sets were used to allow the UAV to explore and map features while always staying in range of a renewal set to satisfy the budget constraint.
Future work could include extending our method of backup trajectory construction to time-varying safe sets, e.g., avoiding a dynamic obstacle, and time-varying budget renewal sets, e.g., a mobile charging station. Also of interest would be to incorporate multiple budget constraints into ReRoot and form a set of Pareto optimal backup trajectories. The active budget constraint in the gatekeeper iteration could then determine which backup to take.